x+2/3=7/8
-9/7 x=27
-17+5/3 x=8
-3(2x-1)=27-2x
7/3-5/2 x=8+2x

Answer by rfer(16322)

1)
x=21/24-16/24
x=5/24
4)
-3(2x-1)=27-2x
-6x+3=27-2x
-4x=24
x=-4